How many positive integers less than 6000 may be written down without using the digits 6,7,8,9 and 0?
Thank you so much

Answer by Edwin McCravy(20054)

We can only use the digits 1,2,3,4, and 5

Number of 1-digit positive integers:

Pick the digit any 5 ways

That's 5 1-digit positive integers

---------------------

Number of 2-digit positive integers:

Pick the first digit any of 5 ways.
Pick the second digit any of 5 ways.

That's 5×5 or 52 2-digit positive integers.

---------------------

Number of 3-digit positive integers:

Pick the first digit any of 5 ways.
Pick the second digit any of 5 ways.
Pick the third digit any of 5 ways

That's 5×5×5 or 53 3-digit positive integers.

---------------------

Number of 4-digit positive integers:

Pick the first digit any of 5 ways.
Pick the second digit any of 5 ways.
Pick the third digit any of 5 ways.
Pick the fourth digit any of 5 ways

That's 5×5×5×5 or 54 4-digit positive integers.

---------------------

Total = 5+52+53+54 = 5+25+125+625=780

Edwin